# Modelling the evolution and influence of dust in cosmological simulations that include the cold phase of the interstellar medium

### v1

@@ 3.7 Clumping factors
As with gas physics in the ISM, dust evolution is a highly multi-scalar process, ranging from dense molecular clouds to the diffuse CGM. In particular, the dense cores of molecular clouds provide sites for the growth of grains through frequent collisions in relatively cold gas (Tacconi et al., 2020). The core densities of molecular clouds typically range from $n_{\rm H}=10^{4}\;{\rm cm}^{-3}$ up to proto-stellar densities. The spatial scales associated with the densest molecular gas is below what we resolve in our cosmological volume simulations. Inability to represent these denser structures skews the densities experienced by gas to lower values and inhibits the accelerated dust evolution inferred for observed molecular clouds. To represent the presence and structure of unresolved collapsed clouds, we experiment with a ‘clumping factor’, $C$. This factor can boost input densities such that the effective gas densities (effective hydrogen number density), $\rho^{\prime}$ ($n^{\prime}_{\rm H}$), used by dense-gas processes (such as accretion and coagulation) are related to the physical densities in our simulations, as
@@
For the noC and maxC10 runs, the clumping factor is lower than or equal to that of the Fiducial run at all densities. We see that for maxC10, the transition to higher $\mathcal{DTZ}$ is pushed to higher densities by $\approx 1$ dex, while grain sizes are systematically larger for $\log_{10}(n_{\rm H}/{\rm cm^{-3}})\lesssim 1.5$ (bottom panel). We note that the maxC10 model does not reach the saturated regime, with an approximately constant $\mathcal{DTZ}$ ratio of $\approx 40\%$ at high ISM densities, as seen in Fiducial. For noC, we see a more extreme shift in the $\mathcal{DTZ}$ transition to higher densities. Interestingly, for $\log_{10}(n_{\rm H}/{\rm cm^{-3}})\gtrsim 1.5$ the noC model yields smaller dust grains (dashed lines in the bottom panel, right axis). These trends in the grain size distribution are indicative of the balance of large-grain dominated populations at injection, shattering at intermediate densities and coagulation in dense gas, alongside the higher rates of gas-phase accretion per unit mass for smaller grains.
@@ 3.8 Depletion
Figure 4: Elemental depletion and total metal depletion of dust constituents for our model. We compare the Fiducial run (blue $\times$ markers), using graphite and olivine grain chemical species, to the grain chemistry variation SilPyro (orange + markers), with graphite and pyroxene chemistry. For comparison, we plot the Milky Way ISM values as empty grey squares (Jenkins 2009, corrected for our assumed $Z_{\odot}=0.0134$), as well as values for a factor 2 reduction in $\delta_{\rm C}$ (Sofia et al., 2011; Parvathi et al., 2012). C is depleted into the homonuclear graphite grains, while O, Mg, Si and Fe are depleted into the heteronuclear silicates. For visibility, the $\delta_{\rm Fe,\,MW}$ is plotted as an upper limit, given the value of -2.2. We also plot the total metal depletion $\delta_{Z}$. We see that the Fiducial run shows best MW-depletion agreement compared to PyroSil, with the exception of Si, which is comparable.
@@
such that more negative values are indicative of larger fractions of X being depleted onto dust grains. Strong depletion of elements is key to limiting grain growth in the ISM of galaxies, and can contribute to a saturation in the $\mathcal{DTZ}$ ratio and thus a near-constant $\mathcal{DTZ}$ ratio measured in massive galaxies (e.g. Zafar & Watson, 2013). While a secondary consideration to the total dust content, elemental depletion helps constrain our model.
We illustrate depletion patterns in Fig. 4 in order to compare how different elements bottleneck grain formation in different models, and how these compare to what we see in the dense ISM of the Milky Way (from Jenkins 2009, corrected to our solar abundance as in Ploeckinger & Schaye 2020). We aggregate gas of $n_{\rm H}>50\;{\rm cm^{-3}}$, to sample the depletion-limited regime as seen in the bottom panel of Fig. 3.
@@ 3.8.1 Silicate grain depletion
In the case of heteronuclear silicate grains, the depletion pattern of constituent elements (O, Mg Si and Fe) depends on the grain chemistry, and its interplay with the yields and diffuse abundance of those elements in the ISM. As such, depletion patterns can provide motivation for our choice of effective grain molecule. We consider both the pyroxene and our fiducial olivine groups as candidate silicate molecules, with the form (Mg,Fe)2Si2O6 and (Mg,Fe)2SiO4, respectively.
@@
While O is observed to be the least depleted silicate constituent, it can still contribute significantly to the silicate mass due to its higher abundance compared to the other constituent elements. We find that using higher olivine and pyroxene silicates both slightly under-depletes O, with the olivine molecule faring slightly better. This is intriguing; olivines represent the most oxygen-rich viable silicate candidates, so a pure olivine model still under-depleting O could be indicative of a potential tension between the metallicities and depletions, perhaps via the metallicity calibration examined further in Appendix C.
@@
We note that tracking more grain chemistry species (e.g. metal oxides, silicon carbide grains), may help to balance the depletion of elements, though this would complicate our model [^13] For example, having multiple grain types containing the same elements requires some decision as to how this depletion is split. and contribute to its memory overhead.
@@ 3.8.2 Graphite grain depletion
In observed molecular clouds, the majority of gas-phase carbon is in CO (Fuente et al., 2019). The triple bond in CO is the strongest molecular bond and CO only solidifies at temperatures $T\lessapprox 70~{}{\rm K}$. We take these properties together to mean that carbon in the CO reservoir, which our simulations do not resolve and which we do not track in our coupled simulations, is not readily available for depletion onto solid grains. We therefore limit carbon depletion by making some fraction of the diffuse-phase element unavailable for depletion. We choose a maximum dust depletion factor of 3 (66% of carbon in dust), following the empirical study of local molecular clouds of Fuente et al. (2019). We note that this limit is only relevant for carbon, and oxygen depletion is instead bottle-necked by the availability of silicate constituent elements.
@@ 4 Results
Here, we present results from our simulations (Table 2), and compare to observations. In §4.1, we first present phase diagrams shaded according to $\mathcal{DTZ}$, isolating the role of different evolutionary processes in the proliferation of dust. We then consider the evolution of cosmic dust density with redshift in §4.2, comparing our Fiducial run to observations, alongside other variation runs isolating different effects. In §4.3-4.6, we then compare galaxy dust relations from our cosmological volume simulations at $z=0$ with low-redshift observational data, focussing on our Fiducial model. Finally, we present and discuss resolved dust maps of a few Fiducial galaxies.
@@ 4.1 Isolating evolutionary processes
In order to show the influence of different dust evolution processes, we compare runs where various processes are isolated. In particular, the SeedOnly run has no ISM dust evolution, only injecting grains from stellar channels. The NoDestruction run then turns on accretion and size transfer processes, leaving processes converting dust mass back into diffuse metals off. These runs are compared to Fiducial in Fig. 5, where we show gas phase diagrams ($T$ as a function of $n_{\rm H}$), the left panel shading by total $\mathcal{DTZ}$ ratio, and centre and right panels shading by the difference in $\mathcal{DTZ}$ with respect to the Fiducial run of SeedOnly and NoDestruction, respectively.
@@ 4.2 Cosmic dust density evolution
For these runs we also show uncoupled equivalents of runs (i.e. where dust model has no influence on other gas physics, see §3.6), indicated by dashed lines, and excepting the SeedOnly run which is only run in an uncoupled context. In other cases we see that the uncoupled runs yield a factor $\sim 2$ lower $\rho_{\rm dust}$ values by $z=1$, as the self-consistent dust-treatment leads to greater grain growth by accretion.
@@ 4.3 Galaxy dust-mass function
In Fig. 7 we plot the Fiducial galaxy-dust mass function (GDMF), the cosmic number density of galaxies per unit logarithmic galaxy dust mass bin width (${\rm dex}^{-1}$), as a function of $M_{\rm dust}$, corrected to our assumed cosmology. To indicate the influence of resolution, we plot the mass function of galaxies with stellar masses $\log_{10}\,M_{\star}/{\rm M_{\odot}}>9$ (i.e. resolved by $\gtrsim 500$ star particles) , as well as for a broader selection, $\log_{10}\,M_{\star}/{\rm M_{\odot}}>8$, using thick and thin blue lines respectively. The narrower selection is plotted down to galaxy masses where it is $>85\%$ complete.
@@ 4.4 Galaxy dust-mass scaling relations
Figure 8: Galaxy dust scaling relations, showing the dust mass ($M_{\rm dust}$) as a function of other baryonic (star and gas) properties of galaxies. In each panel, we show the median relation for galaxies taken from the Fiducial simulation (solid blue line), with shading indicating the 16th-84th percentile range (translucent blue). Individual galaxies are also plotted as low opacity blue data points. The top left panel shows the $M_{\rm\star}$-$M_{\rm dust}$ relation, with the top right panel shows the ${\rm SFR}$-$M_{\rm dust}$ relation, comparing both to DustPedia data presented Bianchi et al. (2018). We also compare $M_{\rm\star}$-$M_{\rm dust}$ specifically to the survey compilation of De Looze et al. (2020). The bottom left panel shows the $M_{\rm neut}$-$M_{\rm dust}$ relation, while the bottom right panel shows $M_{\rm H_{2}}$-$M_{\rm dust}$, comparing to data using the masses of hydrogen phases in DustPedia galaxies, as presented by De Vis et al. (2019), as well as a power-law fit to the Planck-2MRS data (Orellana et al., 2017). To show if the tight $M_{\rm H_{2}}$-$M_{\rm dust}$ relation is driven by the dust coupling, we additionally plot medians in the bottom-right panel and scatter from the FidUncoupled runs, aggregating the P25 dust properties with SOAP.
@@
For the $M_{\star}$-$M_{\rm dust}$ relation, we see a generally concave increasing relation, super-linear for $\log_{10}M_{\star}/{\rm M_{\odot}}<8$, close to linear for $108<\log_{10}M_{\star}/{\rm M_{\odot}}<$ 10, and exhibiting a turnover for $\log_{10}M_{\star}/{\rm M_{\odot}}>10$. This relation is relatively tight, with a scatter of $\lesssim 0.5$ dex at $\log_{10}M_{\star}/{\rm M_{\odot}}=9$. At higher $M_{\star}$, the turnover is induced by an increasing scatter to low $M_{\rm dust}$ in $\log_{10}M_{\star}/{\rm M_{\odot}}=10$ galaxies, increasing to $\sim 2$ dex. Comparing to the observations of Bianchi et al. (2018), we see that the simulated galaxies are consistent with the data, tracing well the upper ridge line amongst the data points. These observations similarly exhibit a marked increase in scatter at higher masses ($M_{\star}/{\rm M_{\odot}}>10$), though they appear to show a larger number of galaxies scattered to low $M_{\rm dust}$ extending to low-intermediate stellar masses ($\log_{10}M_{\star}/{\rm M_{\odot}}>8.5$) [^14] We note that the COLIBRE hybrid AGN feedback model, calibrated in larger volumes than presented here, enhances the scatter in this relation.. We also compare to the multi-survey compilation of De Looze et al. (2020), which exhibits less scatter at intermediate masses, agreeing very well with the Fiducial galaxies in the $8.5\lesssim\log_{10}M_{\star}/{\rm M_{\odot}}\lesssim 10$ range. This compilation also exhibits increased scatter at high $M_{\star}$, though significantly less than seen in the Bianchi et al. (2018) data points, owing to the differing selection functions of the surveys. The Fiducial simulation appears intermediate between these two samples.
@@
The median $M_{\rm dust}$-SFR relation of Fig. 8 (top right) is close to linear in the range $-2<\log_{10}{\rm SFR}/({\rm M_{\odot}\;yr^{-1}})<0.5$, with increasing scatter towards lower SFR values, $\lesssim 0.5$ dex at $1\;{\rm M_{\rm\odot}\;yr^{-1}}$, and reaching $\sim 2$ dex for $0.001\;{\rm M_{\rm\odot}\;yr^{-1}}$. This relation follows the Bianchi et al. (2018) data, albeit with a slightly steeper slope; on the lower side at $-3<\log_{10}{\rm SFR}/({\rm M_{\odot}\;yr^{-1}})<-2$, and on the higher side by $\log_{10}{\rm SFR}/({\rm M_{\odot}\;yr^{-1}})\approx 0$. The observations show a similar trend for the scatter, with slightly larger amplitude. Additional scatter in the observations could be attributed to measurement errors and, at low SFR, the observed $M_{\rm dust}$ distribution being truncated by the lower observational detection limit.
@@
Comparing to gas masses in the bottom row, the $M_{\rm neut}$-$M_{\rm dust}$ relation (bottom left) for median Fiducial galaxies exhibits a steadily increasing, slightly super-linear slope for $\log_{10}(M_{\rm neut}/{\rm M_{\odot}})>8.25$, with a steeper slope below this, owing to a transition to lower dust densities at $\log_{10}(M_{\rm neut}/{\rm M_{\odot}})\approx 8$. This is likely due to a physical transition in galaxies at an ISM density threshold, where the $\mathcal{DTZ}$ relation transitions from being set by the seeding, to being set by a balance between accretion and the availability of diffuse elements for depletion (see following section).

### v2

@@ 3.8 Clumping factors
As with gas physics in the ISM, dust evolution is a highly multi-scalar process, ranging from dense molecular clouds to the diffuse CGM. In particular, the dense cores of molecular clouds provide sites for the growth of grains through frequent collisions in relatively cold gas (Ormel et al., 2009). The core densities of molecular clouds typically range from $n_{\rm H}=10^{4}\;{\rm cm}^{-3}$ up to proto-stellar densities. The spatial scales associated with the densest molecular gas is below what we resolve in our cosmological volume simulations. Inability to represent these denser structures skews the densities experienced by gas to lower values and inhibits the accelerated dust evolution inferred for observed molecular clouds. To represent the presence and structure of unresolved collapsed clouds, we experiment with a ‘clumping factor’, $C$. This factor can boost input densities such that the effective gas densities (effective hydrogen number density), $\rho^{\prime}$ ($n^{\prime}_{\rm H}$), used by dense-gas processes (such as accretion and coagulation) are related to the physical densities in our simulations, as
@@
For the noC and maxC10 runs, the clumping factor is lower than or equal to that of the Fiducial run at all densities. We see that for maxC10, the transition to higher $\mathcal{DTZ}$ is pushed to higher densities by $\approx 1$ dex, while grain sizes are systematically larger for $\log_{10}(n_{\rm H}/{\rm cm^{-3}})\lesssim 1.5$ (bottom panel). We note that the maxC10 model does not reach the saturated regime, with an approximately constant $\mathcal{DTZ}$ ratio of $\approx 40\%$ at high ISM densities, as seen in Fiducial. For noC, we see a more extreme shift in the $\mathcal{DTZ}$ transition to higher densities. Interestingly, for $\log_{10}(n_{\rm H}/{\rm cm^{-3}})\gtrsim 1.5$ the noC model yields smaller dust grains (dashed lines in the bottom panel, right axis). These trends in the grain size distribution are indicative of the lasting imprint of the grain sizes assumed for seed grains in low density gas where dust evolution proceeds slowly, i.e. reflective of seed grains having 90% in large grains by mass (§ 3.2.4), followed by efficient shattering at intermediate densities and coagulation in dense gas, balanced with the higher rates of gas-phase accretion per unit mass for smaller grains.
@@ 3.9 Depletion
Figure 4: Elemental depletion and total metal depletion of dust constituents for our model. We compare the Fiducial run (blue $\times$ markers), using graphite and olivine grain chemical species, to the grain chemistry variation SilPyro (orange + markers), with graphite and pyroxene chemistry. For comparison, we plot the Milky Way ISM values as empty grey squares (Jenkins 2009, corrected for our assumed $Z_{\odot}=0.0134$), as well as values for a factor 2 reduction in $\delta_{\rm C}$ (Sofia et al., 2011; Parvathi et al., 2012). C is depleted into the homonuclear graphite grains, while O, Mg, Si and Fe are depleted into the heteronuclear silicates. For visibility, the $\delta_{\rm Fe,\,MW}$ is plotted as an upper limit, given the value of -$2.$ We also plot the total metal depletion $\delta_{Z}$. We see that the Fiducial run shows best MW-depletion agreement compared to PyroSil, with the exception of Si, which is comparable.
@@
such that more negative values are indicative of larger fractions of X being depleted onto dust grains. Strong depletion of elements is key to limiting grain growth in the ISM of galaxies, and can contribute to a saturation in the $\mathcal{DTZ}$ ratio and thus a near-constant $\mathcal{DTZ}$ ratio measured in massive galaxies (e.g. Zafar & Watson, 2013; Rémy-Ruyer et al., 2014; De Vis et al., 2019). While a secondary consideration to the total dust content, elemental depletion helps constrain our model.
We illustrate depletion patterns in Fig. 4 in order to compare how different elements bottleneck grain formation in different models, and how these compare to what we see in the dense ISM of the Milky Way (from Jenkins 2009, assuming the strong depletion case, $F_{\star}=1$, corrected to our solar abundance as in Ploeckinger & Schaye 2020). We aggregate gas of $n_{\rm H}>50\;{\rm cm^{-3}}$, to sample the depletion-limited regime as seen in the bottom panel of Fig. 3.
@@ 3.9.1 Silicate grain depletion
In the case of heteronuclear silicate grains, the depletion pattern of constituent elements (O, Mg Si and Fe) depends on the grain chemistry, and its interplay with the yields and gas-phase abundance of those elements in the ISM. As such, depletion patterns can provide motivation for our choice of effective grain molecule. We consider both the pyroxene and our fiducial olivine groups as candidate silicate molecules, with the form (Mg,Fe)2Si2O6 and (Mg,Fe)2SiO4, respectively.
@@
While O is observed to be the least depleted silicate constituent, it can still contribute significantly to the silicate mass due to its higher abundance compared to the other constituent elements. We find that using higher olivine and pyroxene silicates both slightly under-depletes O, with the olivine molecule faring slightly better. This is intriguing; olivines represent the most oxygen-rich viable silicate candidates, so a pure olivine model still under-depleting O could be indicative of a potential tension between the metallicities and depletions, perhaps via the metallicity calibration examined further in Appendix C. This difficulty in explaining O depletions has been noted previously (e.g. Whittet, 2010).
@@
We note that laboratory studies comparing to observations in the X-ray can provide evidence for certain (olivine or pyroxene) silicates or some relative mix between the two (see e.g. Zeegers et al., 2019; Rogantini et al., 2020; Psaradaki et al., 2023). While a single silicate group is clearly a simplification, and tracking more grain chemistry species (e.g. including metal oxides, silicon carbide grains), may help to balance the depletion of elements, this would complicate our model [^15] For example, having multiple grain types containing the same elements requires some decision as to how this depletion is split. and contribute to its memory overhead. We therefore use olivine as a viable, effective molecule that we find to reasonably reproduce depletion through its fixed ratio of constituent elements is therefore our preference.
@@ 3.9.2 Graphite grain depletion
In observed molecular clouds, the majority of gas-phase carbon is in CO (Fuente et al., 2019). The triple bond in CO is the strongest molecular bond and CO only solidifies at temperatures $T\lessapprox 70~{\rm K}$. We take these properties together to mean that carbon in the CO reservoir, which our simulations do not resolve and which we do not track in our coupled simulations, is not readily available for depletion onto solid grains. We therefore limit carbon depletion by making some fraction of the gas-phase element unavailable for depletion. We choose a maximum dust depletion factor of 3 (66% of carbon in dust), following the empirical study of local molecular clouds of Fuente et al. (2019). We note that this limit is only relevant for carbon, and oxygen depletion is instead bottle-necked by the availability of silicate constituent elements.
@@
3.10 Resolution
While we focus on the m6 resolution in this work, with gas (DM) particles of $1.84\times 10^{6}\;{\rm M_{\odot}}$ ($2.42\times 10^{6}\;{\rm M_{\odot}}$) , the dust model presented here is intended to be used across a range of resolutions, and as a single component of the multi-resolution COLIBRE simulation suite. In particular, we defer plots showing the convergence properties of dust (particularly the $M_{\star}$-$M_{\rm dust}$ relation and grain size ratio as a function of $M_{\star}$) to the presentation of the COLIBRE simulations as a whole by Schaye et al. (2025) in context with the convergence of other aspects of galaxies. With COLIBRE we target ‘weak convergence’ (e.g. Schaye et al., 2015), aiming to achieve a consistent repoduction of target datasets through the choice and calibration of key parameters. This calibration process happens in concert with the other modules that constitute COLIBRE, with feedback calibration detailed in Chaikin et al. (2025).
Elements of the modelling were designed to support a range of resolutions, for example by the implementation of a clumping factor (§ 3.8) intended to represent unresolved structures like molecular clouds and their dense cores at intermediate and low resolutions, as opposed to an explicit sub-grid model of molecular clouds in simulation. The dust properties in COLIBRE are generally converged for galaxies resolved by $\gtrsim$1000 star particles, or $\gtrsim 2\times 10^{9}\;{\rm M_{\odot}}$ at m6 resolution. At the highest masses, $\gtrapprox 10^{10.7}\;{\rm M_{\odot}}$, convergence deteriorates somewhat owing to the role of AGN feedback calibration. Dust convergence properties are discussed in more detail in § 7.2.5 of Schaye et al. (2025).
@@ 4 Results
Here, we present results from our simulations (Table 1), and compare to observations. In §4.1 we first present and discuss some resolved dust maps for some of our Fiducial run galaxies. In §4.2, we present phase diagrams shaded according to $\mathcal{DTZ}$, isolating the role of different evolutionary processes in the proliferation of dust. We then consider the evolution of cosmic dust density with redshift in §4.3, comparing our Fiducial run to observations, alongside other variation runs isolating different effects. In §4.4-4.7, we then compare galaxy dust relations from our cosmological volume simulations at $z=0$ with low-redshift observational data, focussing on our Fiducial model.
4.1 Resolved properties
[…]
First, we briefly present a gallery of randomly oriented (i.e. box-projected) dust maps for a handful of galaxies taken from our Fiducial model in Fig. 6. These are the 6 highest stellar mass galaxies, defined within a 50 ckpc spherical aperture, that also have $10\%$ of their baryonic mass constituted by gas. For each galaxy, we map the highest dust surface density ($\Sigma_{\rm d}$) pixels with colour shading, using a kernel density projection of the SPH gas particles, via the swiftgalaxy161616github.com/SWIFTSIM/swiftgalaxy tool. These are plotted alongside neutral and molecular hydrogen contours to show the relatively cold-dense gas regions. Colour shadings illustrate various properties of the dust population; the local $\mathcal{DTZ}$ ratio, the small-to-large grain ratio, the ratio of silicate to graphite/carbon grains and the ratio of forsterite (Mg-endmember) to fayalite (Fe-endmember) silicates.
@@ 4.2 Isolating evolutionary processes
In order to show the influence of different dust evolution processes, we compare runs where various processes are isolated. In particular, the SeedOnly run has no ISM dust evolution, only injecting grains from stellar channels. The NoDestruction run then turns on accretion and size transfer processes, leaving processes converting dust mass back into gas-phase metals off. These runs are compared to Fiducial in Fig. 5, where we show gas phase diagrams ($T$ as a function of $n_{\rm H}$), the left panel shading by total $\mathcal{DTZ}$ ratio, and centre and right panels shading by the difference in $\mathcal{DTZ}$ with respect to the Fiducial run of SeedOnly and NoDestruction, respectively.
@@ 4.3 Cosmic dust density evolution
For these runs we also show uncoupled equivalents of runs (i.e. where dust model has no influence on other gas physics, see §3.7), indicated by dashed lines, and excepting the SeedOnly run which is only run in an uncoupled context. In other cases we see that the uncoupled runs yield a factor $\sim 2$ lower $\rho_{\rm dust}$ values by $z=1$, as the self-consistent dust-treatment leads to greater grain growth by accretion. In addition, for the SeedOnly run we also show an orange dotted curve illustrating $\rho_{\rm dust}$ boosted by an order of magnitude, and empty orange circles showing the same boost for only dust within the $M_{\star}>10^{9}\;{\rm M_{\odot}}$ sample. This is intended as a rough approximation of dust yield levels comparable to e.g. Dwek (1998). We see that a high-yield paradigm appears better at reproducing the observations in lieu of evolutionary processes. However, we see that only counting dust within $M_{\star}>10^{9}\;{\rm M_{\odot}}$ galaxies is still significantly low relative to the data so would require dust growth. This shows that adopting higher dust yields (discussed in § 3.2) could be a means to reproduce observations regulated with milder evolutionary processes (e.g. a lower clumping factor, § 3.8), as found by e.g. Choban et al. (2025).
@@ 4.4 Galaxy dust-mass function
In Fig. 8 we plot the Fiducial galaxy-dust mass function (GDMF), the cosmic number density of galaxies per unit logarithmic galaxy dust mass bin width (${\rm dex}^{-1}$), as a function of $M_{\rm dust}$, corrected to our assumed cosmology. To indicate the influence of resolution, we plot the mass function of galaxies with stellar masses $\log_{10}\,M_{\star}/{\rm M_{\odot}}>9$ (i.e. resolved by $\gtrsim 500$ star particles) , as well as for a broader selection, $\log_{10}\,M_{\star}/{\rm M_{\odot}}>8$, using thick and thin blue lines respectively. The narrower selection is plotted down to galaxy masses where it is $85\%$ complete.
@@ 4.5 Galaxy dust-mass scaling relations
Figure 9: Galaxy dust scaling relations, showing the dust mass ($M_{\rm dust}$) as a function of other baryonic (star and gas) properties of galaxies. In each panel, we show the median relation for galaxies taken from the Fiducial simulation (solid blue line), with shading indicating the 16th-84th percentile range (translucent blue). Individual galaxies are also plotted as low opacity blue data points. The top left panel shows the $M_{\rm\star}$-$M_{\rm dust}$ relation, with the top right panel shows the $\rm SFR$$M_{\rm dust}$ relation, comparing both to DustPedia data presented Bianchi et al. (2018). We also compare $M_{\rm\star}$-$M_{\rm dust}$ specifically to the survey compilation of De Looze et al. (2020) (indicated for each contributing survey using grey markers). The bottom left panel shows the $M_{\rm neut}$-$M_{\rm dust}$ relation, while the bottom right panel shows $M_{\rm H_{2}}$-$M_{\rm dust}$, comparing to data using the masses of hydrogen phases in DustPedia galaxies, as presented by De Vis et al. (2019), as well as a power-law fit to the Planck-2MRS data (Orellana et al., 2017). To show if the tight $M_{\rm H_{2}}$-$M_{\rm dust}$ relation is driven by the dust coupling, we additionally plot medians in the bottom-right panel and scatter from the FidUncoupled runs, aggregating the P25 dust properties with SOAP.
@@
For the $M_{\star}$-$M_{\rm dust}$ relation, we see a generally concave increasing relation, super-linear for $\log_{10}M_{\star}/{\rm M_{\odot}}<8$, close to linear for $8<\log_{10}M_{\star}/{\rm M_{\odot}}<10$, and exhibiting a turnover for $\log_{10}M_{\star}/{\rm M_{\odot}}>10$. This relation is relatively tight, with a scatter of $\lesssim 0.5$ dex at $\log_{10}M_{\star}/{\rm M_{\odot}}=9$. At higher $M_{\star}$, the turnover is induced by an increasing scatter to low $M_{\rm dust}$ in $\log_{10}M_{\star}/{\rm M_{\odot}}=10$ galaxies, increasing to $\sim 2$ dex. Physically, this scatter to lower $M_{\rm dust}$ owes to increasing passive fractions at high $M_{\star}$, as evidenced by the relatively tight relation displayed by the $M_{\rm dust}$-SFR relation (top right, discussed below), tightening with $M_{\star}$. Comparing to the observations of Bianchi et al. (2018), we see that the simulated galaxies are consistent with the data, tracing well the upper ridge line amongst the data points. These observations similarly exhibit a marked increase in scatter at higher masses ($M_{\star}/{\rm M_{\odot}}>10$), though they appear to show a larger number of galaxies scattered to low $M_{\rm dust}$ extending to low-intermediate stellar masses ($\log_{10}M_{\star}/{\rm M_{\odot}}>8.5$) [^17] We note that the COLIBRE hybrid AGN feedback model, calibrated in larger volumes than presented here, enhances the high-mass scatter in this relation by 0.2-0.3 dex for galaxies of $\log_{10}M_{\star}/{\rm M_{\odot}}>10$ in a 1003 Mpc3 volume at m6 resolution.. We also compare to the multi-survey compilation of De Looze et al. (2020), which exhibits less scatter at intermediate masses, agreeing very well with the Fiducial galaxies in the $8.5\lesssim\log_{10}M_{\star}/{\rm M_{\odot}}\lesssim 10$ range. This compilation also exhibits increased scatter at high $M_{\star}$, though significantly less than seen in the Bianchi et al. (2018) data points. This owes to the differing selection functions of the surveys, in particular the heavy sampling of Virgo and Fornax clusters given the median sample distance of 20 Mpc, and the stronger representation of early type galaxies in DustPedia scattering to lower dust-to-stellar ratios (Davies et al., 2019). The Fiducial simulation appears intermediate between these two samples.
@@
The median $M_{\rm dust}$-SFR relation of Fig. 9 (top right) is close to linear in the range $2<\log_{10}{\rm SFR}/({\rm M_{\odot}\;yr^{-1}})<0.5$, with increasing scatter towards lower SFR values, $\lesssim 0.5$ dex at $1\;{\rm M_{\rm\odot}\;yr^{-1}}$, and reaching $\sim 2$ dex for $0.001\;{\rm M_{\rm\odot}\;yr^{-1}}$. This relation follows the Bianchi et al. (2018) data, albeit with a slightly steeper slope; on the lower side at $3<\log_{10}{\rm SFR}/({\rm M_{\odot}\;yr^{-1}})<-2$, and on the higher side by $\log_{10}{\rm SFR}/({\rm M_{\odot}\;yr^{-1}})\approx 0$. The observations show a similar trend for the scatter, with slightly larger amplitude. Additional scatter in the observations could be attributed to the larger uncertainties in measuring dim FIR sources corresponding to the lowest dust masses.
@@
Comparing to gas masses in the bottom row, the $M_{\rm neut}$-$M_{\rm dust}$ relation (bottom left) for median Fiducial galaxies exhibits a steadily increasing, slightly super-linear slope for $\log_{10}(M_{\rm neut}/{\rm M_{\odot}})>8.25$, with a steeper slope below this, owing to a transition to lower dust densities at $\log_{10}(M_{\rm neut}/{\rm M_{\odot}})\approx 8$. This is likely due to a physical transition in galaxies at an ISM density threshold, where the $\mathcal{DTZ}$ relation transitions from being set by the seeding, to being set by a balance between accretion and the availability of gas-phase elements for depletion (see following section).